OpenStudy (anonymous):

What is the slope of the line 4x + 6y = 12? please help me friends (:

OpenStudy (anonymous):

Someone already tried to help you. Put it into the form y = mx + b

OpenStudy (anonymous):

please help me @SACAPUNTAS

OpenStudy (anonymous):

Algebra. Do the same thing on each side of the equal sign until it's in a form you want. For example, if you subtract 4x from one side, you have to subtract 4x from the other side, too. \[4x+6y=12\]\[\cancel{4x-4x}+6y=12-4x\]

OpenStudy (vjordillo):

\[m=\frac{ -A }{ B } \] in \[Ax ^{2} + Bx +C = 0\]

OpenStudy (anonymous):

A. 2/3 B. 2 C. -2/3 D.6 these are the answer choices please choosethe right one (: thanks (:

OpenStudy (anonymous):

I showed you how to get the answer. I'm not going to just tell you the answer.